मराठी

The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1.

Advertisements
Advertisements

प्रश्न

The probability that a student will pass the final examination in both English and Hindi is 0.5 and the probability of passing neither is 0.1. If the probability of passing the English Examination is 0.75. What is the probability of passing the Hindi Examination?

बेरीज
Advertisements

उत्तर

Let A and B be the events of passing English and Hindi examinations, respectively.

Accordingly, we have:

P(A and B) = 0.5

 P(not A and not B) = 0.1 [i.e. P(A' ∩ B') = 0.1]

P(A) = 0.75

Now, P(A∪B) + P(A' ∩ B') = 1

⇒ P(A∪B) = 1 - P(A' ∩ B')

                  = 1 -0.1 = 0.9

By addition theorem, we have:

P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

⇒ 0.9 = 0.75 + P (B)  - 0.5

⇒ P(B) = 0.9 - 0.75 + 0.5

⇒ P(B) = 0.65

Thus, the probability of passing the Hindi examination is 0.65.

shaalaa.com
Concept of Probability - Probability of 'Not', 'And' and 'Or' Events
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 14: Probability - EXERCISE 14.2 [पृष्ठ ३०७]

APPEARS IN

एनसीईआरटी Mathematics [English] Class 11
पाठ 14 Probability
EXERCISE 14.2 | Q 20. | पृष्ठ ३०७
आर.डी. शर्मा Mathematics [English] Class 11
पाठ 33 Probability
Exercise 33.4 | Q 13 | पृष्ठ ६८

संबंधित प्रश्‍न

If `2/11` is the probability of an event, what is the probability of the event ‘not A’.


A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P(A or B).


In Class XI of a school 40% of the students study Mathematics and 30% study Biology. 10% of the class study both Mathematics and Biology. If a student is selected at random from the class, find the probability that he will be studying Mathematics or Biology.


In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing at least one of them is 0.95. What is the probability of passing both?


In a certain lottery, 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy one ticket.


In a simultaneous throw of a pair of dice, find the probability of getting a doublet


In a simultaneous throw of a pair of dice, find the probability of getting a doublet of prime numbers


In a simultaneous throw of a pair of dice, find the probability of getting a sum greater than 9


In a simultaneous throw of a pair of dice, find the probability of getting an even number on one and a multiple of 3 on the other


In a simultaneous throw of a pair of dice, find the probability of getting a sum less than 7


In a simultaneous throw of a pair of dice, find the probability of getting a sum more than 7


In a simultaneous throw of a pair of dice, find the probability of getting neither a doublet nor a total of 10


In a simultaneous throw of a pair of dice, find the probability of getting a total of 9 or 11


Three coins are tossed together. Find the probability of getting at least two heads


Three coins are tossed together. Find the probability of getting at least one head and one tail.

 

Two dice are thrown. Find the odds in favour of getting the sum  What are the odds against getting the sum 6?


What are the odds in favour of getting a spade if the card drawn from a well-shuffled deck of cards? What are the odds in favour of getting a king?

 

A box contains 10 red marbles, 20 blue marbles and 30 green marbles. 5 marbles are drawn at random. From the box, what is the probability that all are blue?


A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is red or even numbered.


Fill in the blank in the table:

P (A) P (B) P (A ∩ B) P(A∪ B)
\[\frac{1}{3}\] \[\frac{1}{5}\] \[\frac{1}{15}\] ......

If A and B are two events associated with a random experiment such that P(A) = 0.3, P (B) = 0.4 and P (A ∪ B) = 0.5, find P (A ∩ B).


If A and B are two events associated with a random experiment such that
P(A) = 0.5, P(B) = 0.3 and P (A ∩ B) = 0.2, find P (A ∪ B).


One of the two events must happen. Given that the chance of one is two-third of the other, find the odds in favour of the other.


A card is drawn from a deck of 52 cards. Find the probability of getting an ace or a spade card.


In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing at least one of them is 0.95. What is the probability of passing both?


The probability that a leap year will have 53 Fridays or 53 Saturdays is


A and B are two events such that P (A) = 0.25 and P (B) = 0.50. The probability of both happening together is 0.14. The probability of both A and B not happening is


Three integers are chosen at random from the first 20 integers. The probability that their product is even is


A bag contains 5 black balls, 4 white balls and 3 red balls. If a ball is selected randomwise, the probability that it is black or red ball is


Three numbers are chosen from 1 to 20. The probability that they are not consecutive is


In a certain lottery 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy 10 tickets.


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×